OpenStudy (anonymous):

What is the x-intercept of the line with the equation? 1/3x + y = -27

OpenStudy (anonymous):

in x-intercept, y = 0. So put y = 0 in the equation and find the answer

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@EinsteinPart2

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }x + y = -27\] Put y = 0 \[\frac{ 1 }{ 3 }x + 0 = -27\] \[x = -27 * 3\]

OpenStudy (anonymous):

@sangya21 Is accurate. \[x = -27 × 3\]

OpenStudy (anonymous):

Just be sure not to forget to substitute Y for zero \[Y \sqrt{O}\]